Determine wheather the given ordered pair is a solution of the given equation.
2x-5y=18;(4,2)

Answer by nerdybill(7384)


2x-5y=18;(4,2)
.
An "ordered pair" is denoted as (x,y)
So, in this problem, they are asking if
x were equal to 4
and
y were equal to 2
is that a solution for:
2x-5y=18
.
To check,
Plug the given values back into the equation and check:
2x-5y=18
2(4)-5(2)=18
8-10=18
-2 = 18 (NOT true)
.
Therefore (4,2) is NOT a solution.